# Inverse design of potential metastructures inspired from Indian medieval architectural elements

### v2

@@ Results
The comparison of generated metastructure for one test sample from the test data (10% of available 720 samples) is provided in Fig. 8. Figures 8(a), (b) show the test metastructure and its transmissibility obtained by the FEM method which is assumed to be true (input) transmissibility for inverse design. The predicted metastructure by the inverse design model exactly matches the original metastructure with an SSIM score of 0.991 as shown in Figure 8(c). Forward analysis is performed on both the original and the predicted metastructures. The predicted transmissibilities are compared with the input (true) transmissibility in Fig. 8(d) and (e). It can be observed that MSEs for transmissibility obtained from the original metastructure and the predicted metastructure are 0.0834 and 0.1939 respectively. The MSE for predicted transmissibility by forward analysis on the predicted metastructure is comparatively larger because of the existence of noises in the predicted metastructure.
Validation on Artificial Transmissibility:
@@
The AI-based inverse design model is further validated for practical deployment to design a metastructure facilitating the formation of a desired frequency bandgap. Hence, an artificial transmissibility is created having the desired bandgap and fed to the inverse design model. The generated metastructure and the corresponding transmissibilities obtained by the forward analysis model and FEM are shown in Fig. 9. The obtained metastructure is the combination of a honeycomb and hierarchical honeycomb substructures. The MSE between the predicted transmissibility and the input transmissibility is 0.7408.
Figure 9: AI-based inverse design of metastructure for a desired bandgap. (a) Artificial (input) transmissibility having the desired bandgap as highlighted. (b) The generated metastructure obtained from the inverse design model, which facilitates the desired bandgap. (c) Validation of the generated metastructure design by comparing the transmissibilities obtained by the forward analysis model and FEM.
Discussion
@@
It has been found experimentally and analytically from the previous study that a simple lattice structure has the ability to block waves for a certain frequency range when it passes longitudinally throughout the lattice. Inspired from the architectural patterns, we have designed mechanical metamaterial lattices with diverse interlaced structures. A comprehensive study has been performed through simulation-based analysis using Ansys. Experiments have been conducted to validate the transmissibility property corresponding to different lattice configurations. This work shows the blueprint of metamaterials, including the fact that the effective properties of the lattice are contingent on collective patterns and subpatterns of the unit cells. Furthermore, the tunability of the bandgaps through design resulted in larger and multiple bandgaps occurring in Frequency Response Function (FRF) plots. The reported band gap caused by the lattice structures can be further modified by adding metallic inserts with different combinations of insert patterns. These metallic inserts interact with the geometry and due to the large density difference as compared to lattice material (PLA), they act as local resonators, providing flexibility to adjust bandgaps. Some significant negative peaks are obtained from the transmissibility plots that signify wave attenuation and show good agreement with the simulation results, as shown in Fig. 5. The discrepancy in some amplitude and peak are due to nonconsideration of damping in simulation. Apart from damping, printing speed, orientation, infill pattern, and infill density of 3D printing are also affected as discussed and shown in Supplementary Table S2. These findings unveil a deep connection between lattices and band gaps and have broad potential applications in the field of vibration isolation for high-speed trains and in serving as surface waveguides and absorbers.
@@
The second part of this study on the inverse design of metastructure using geometric attention-based deep architecture is a benchmark in the field of inverse design of complex structure whether it is metastructure for the desired mechanical response or any other complex pattern for desired physical properties. The major challenge in the training of an AI-based inverse design model is the availability of sufficient number of training samples. The higher the complexity of the structure, the higher the number of samples required for the training. Our proposed model uses FEM to learn the dependencies of adjacent connections in the geometrical structure. Therefore, it can easily extract the information embedded in the complex structure and learn the relationship between the input geometry and the output transmissibility. Further, the multiscale geometrical attention mechanism provides the capability to learn the multiple levels of complexity. With the same concept, multiscale feature transformation using the multiscale residual network enables the inverse design model to reconstruct multiple levels of geometry, like a honeycomb and hierarchical honeycomb substructures. Therefore, the proposed framework of the geometrical attention-based feature learning and the inverse design can be useful for any other scientific design problem.
@@
  • Discovery of novel functionality in a set of ancient patterns from the mechanical metastructure perspective such as vibration transmission and attenuation.
@@
  • Creation of novel AI-based forward analysis model using multiscale geometric attention mechanism which can predict the transmissibility as well as bandgap of a metastructure subjected to base excitation.
[…]
  • Further, we have been able to create complex multiscale patterns successfully and predict their bandgap characteristics through the transmissibility plot using the current AI model. We believe this will facilitate the generation of many customizable complex patterns for mechanical metastructures efficiently.
@@
  • A novel inverse design model of metastructure based on NLP inspired multiheaded attention mechanism with 1D position encoding and multiscale residual network has also been demonstrated by providing artificial transmissibility for a desired frequency bandgap. The inverse design model is found to be much more efficient to reconstruct the complex metastructure with similarity index of 0.991.
Conclusions
@@
This research study highlights the translational application potential of vast range of patterns and motifs developed in the field of architecture into the metamaterial-based wave attenuation. Thus, the utilitarian values of aesthetic constructs in the field of vibration and sound engineering may open up new systems for wave control including bandwidth control, subwavelength acoustic imaging, and generation of acoustic black holes. Additionally, a systematic investigation process has been created on the basis of root lattice identification, multi-scale lattice formation through a hierarchic system, development of hybrid lattice structure, and the effect of interlacing in the vibration mitigation potential initiated from the historic geometric patterns found in the tomb of I’timad-ud-Daulla. In general, the simple lattice structures typically act as attenuators, however, this study demonstrates that the interlaced structure offers an enhanced vibration attenuation property depending on the interlace geometry with improved control over the attenuation frequency range. Such effects can be further enhanced, in the presence of and through strategic positioning of the metallic inserts inside the lattice structure which may result in significant expansion of the bandwidth. This method has been enhanced with the help of AI-based inverse design models to predict metastructures that exhibit the desired bandgap and the corresponding transmissibility. Although this study concentrates on predominantly honeycomb-shaped interlaced structures, the principle of incorporating mass resonators can be applied to other geometries to manipulate bandgap frequencies. Moreover, the concept may be extended to three-dimensional interlaced structures with diverse lattice and interlace configurations across multiple length scales. Another possibility of future expansion is related to the insertion of smart materials in the lattice structure such that the stiffness and the damping can be controlled over time; thus, transforming it into 4-D metamaterial.
Methods
Experimental testing
The study utilizes a non-contact technique to measure vibration to assess the response of different latticed metastructures. Nine 3D printed samples, each showing unique metastructure, are subjected to dynamic testing using a 3D Laser Doppler vibrometer (LDV) from Polytec (Details of the experimental setup presented in Supplementary Fig. S10). These samples are attached to an Electrodynamic shaker system in cantilever mode. The laser head of LDV captures surface velocity by analyzing the phase shift between the incident and reflected laser beams. Data acquisition and signal processing are executed through the NI-DAQ system. The base excitation technique is utilized to determine transmissibility. In this method, the LDS shaker generates 3200 Fast Fourier Transform (FFT) lines, and the pseudo-random excitation with a frequency range of 0 to 10 KHz. Displacements of the top and bottom surfaces of the samples are recorded to calculate transmissibility after post-processing of the data.
Finite element simulations
Finite element analysis-based simulation has been carried out using Ansys, employing the Ansys Parametric Design Language (APDL) solver for all the nine structures, and the host material is chosen as PLA (detailed properties given in Supplementary Table S3). The entire structure was discretized by using tetrahedron elements with a 2 mm element size, to ensure convergence for interlaced geometries. The smaller edges of the base structure were fixed, while the longer edges remained free. A dynamic force of amplitude 10 N was applied to the base, and the response was calculated on the opposite phase. Modal analysis was performed to identify all modes up to 10 KHz frequency, followed by harmonic analysis to obtain the transmissibility. The mode superposition method was used to link the modal analysis results with the harmonic analysis input. Frequency response functions (FRFs) were generated over the frequency range of 0 to 10 KHz with a 10 Hz frequency resolution; this step size is preferable to accurately capture the response. A structural damping ratio of 0.001 was incorporated to account for the viscoelastic effect of PLA substrate.
@@ Acknowledgements
Author information
Authors and Affiliations
Department of Mechanical Engineering, Indian Institute of Technology Kanpur, India
Bishakh Bhattacharya, Tanuj Gupta, Arun Kumar Sharma, Ankur Dwivedi, Subhadeep Sahana, Ashish Awasthi
Department of Precision and Microsystems Engineering, Delft University of Technology, Netherland
Vivek Gupta
Department of Mechanical Engineering, Motilal Nehru National Institute of Technology Allahabad, India
Suryansh Pathak
@@
Bishakh Bhattacharya: Conceived the primary idea, reviewed the manuscript, critically examined the results, checked the data, wrote the manuscript, and acquired funding. Tanuj Gupta: Expanded the idea jointly, modeled all the structures, and designed and performed numerical simulations. Fabricated the samples and carried out related experiments and material testing, documented the results, reviewed the manuscript, critically examined the results, and checked the data, wrote the manuscript. Arun Kumar Sharma: Concived the idea of AI-based inverse design of metastructures, and implemented a new strategy of feature learning using multiscale Geometric Attention and the idea of reproducing the geometry with multiple levels of complexity using the multiscale residual network. Code Implementation, training, and validation of the AI models and writing the AI-based inverse design in the manuscript. Critically examined the manuscript. Ankur Dwivedi: Conceived and expanded the idea jointly in the inverse design of metastructures, critically examined the results, checked the data, wrote and reviewed the manuscript. Vivek Gupta: Expanded the idea jointly, reviewed the manuscript, critically examined the results, checked the data, and wrote the manuscript. Subhadeep Sahana: Carried out numerical simulations to generate synthetic data for training the AI model and critically analyzed the paper for logical conclusions. Suryansh Pathak: Conducted the experiments, performed numerical simulations, and checked the data. Ashish Awasthi: Carried out numerical simulations to generate synthetic data for training the AI model.

### v3

@@ 4.5.1 Validation on test samples of metastructure:
The comparison of generated metastructure for one test sample from the test dataset (10% of available 720 samples) is provided in Fig. 12. Figures 12(a), (b) show the test metastructure and its transmissibility obtained by the FEM method which is assumed to be true (input) transmissibility for inverse design. The predicted metastructure by the inverse design model exactly matches the original metastructure with the similarity score of 0.991 as shown in Figure 12(c). Forward analysis is performed on both the original and the predicted metastructures. The predicted transmissibilities are compared with the input (true) transmissibility in Fig. 12(d) and (e). It can be observed that MSEs for transmissibility obtained from the original metastructure and the predicted metastructure are 0.0834 and 0.1939 respectively. The MSE for predicted transmissibility by the forward analysis on the predicted metastructure is comparatively larger because of the existence of noises in the predicted metastructure.
Figure 13: Additional validation results on multiscale metastructures with different combinations of the unit cells (Honeycomb as a primary unit cell and auxetic, star, hierarchical honeycomb interlacing as secondary and tertiary).
To further demonstrate the performance of the inverse design framework, we have provided comparisons of the generated metastructure with the original metastructure having different combinations of unit cells in Fig. 13. The first column of Fig. 13(a) has honeycomb and star interlacing, Fig. 13(b) has honeycomb, star, and auxetic interlacing, Fig. 13(c) has auxetic, circle, honeycomb, and hierarchical honeycomb interlacing with the similarity score of 0.992, 0.992, and 0.990 respectively. The third column of Fig. 13 shows the comparisons of predicted transmissibility of the generated metastructure by the forward analysis model and the true transmissibility evaluated by the FEM on the original metastructure with the MSE of 0.0669, 0.0763, and 0.0741 in respective cases.
4.5.2 Validation on tailored transmissibility:
@@
The AI-based inverse design model is further validated for practical deployment to design a metastructure facilitating the formation of a desired frequency bandgap. Hence, an artificial transmissibility is created with the desired bandgap and fed to the inverse design model. The generated metastructure and the corresponding transmissibilities obtained by the forward analysis model and FEM analysis are shown in Fig. 14. The first column in Fig. 14 shows the artificial spectrum created for the tailored bandgap, and the second column presents the metastructures generated from the inverse design model for the input transmissibility with tailored bandgap. The third column compares the predicted transmissibility by the forward model and the true transmissibility evaluated by the FEM on the generated metastructure.
Figure 14: AI-based inverse design of metastructure for tailored bandgaps. (a) Transmissibility having tailored bandgap at high frequency. (b) Transmissibility having tailored wider bandgap. (c) Transmissibility having multiple tailored bandgaps ranging from low frequency to high frequency.
The predicted metastructure design is the combination of a honeycomb and hierarchical honeycomb substructures as shown in Fig. 14(a). As the attenuation band is in the higher frequency region, hence the inverse design model generates a hierarchical substructure in the generated metamaterial design as previously shown in Fig. 5(a). The MSE between the predicted transmissibility and the input transmissibility is 0.0958. Figures 14 (b) and (c) have interlaced structures such as honeycomb, auxetic, circle, and star in predicted metastructure design. The tailored transmissibility has a wider bandgap and multiple bandgaps in figures 14 (b) and (c) respectively. Hence the inverse design model generated auxetic substructures in the generated metamaterial design.
5 Discussion
@@
It has been found experimentally and analytically from the previous study that a simple lattice structure has the ability to block waves for a certain frequency range when it passes longitudinally throughout the lattice. Inspired from the architectural patterns, we have designed mechanical metamaterial lattices with diverse interlaced structures. A comprehensive study has been performed through simulation-based analysis using Ansys. Experiments have been conducted to validate the transmissibility property corresponding to different lattice configurations. This work shows the blueprint of metamaterials, including the fact that the effective properties of the lattice are contingent on collective patterns and subpatterns of the unit cells. Furthermore, the tunability of the bandgaps through design resulted in larger and multiple bandgaps occurring in Frequency Response Function (FRF) plots. The reported band gap caused by the lattice structures can be further modified by adding metallic inserts with different combinations of insert patterns. These metallic inserts interact with the geometry and due to the large density difference as compared to lattice material (PLA), they act as local resonators, providing flexibility to adjust bandgaps. Some significant negative peaks are obtained from the transmissibility plots that signify wave attenuation and show good agreement with the simulation results, as shown in Fig. 7. The discrepancy in some amplitudes and peaks is due to the nonconsideration of structural damping in the simulation. Apart from damping, printing speed, orientation, infill pattern, and infill density of 3D printing are also affected. These findings unveil a deep connection between lattices and band gaps and have broad potential applications in the field of vibration isolation for high-speed trains and in serving as surface waveguides and absorbers.
@@
The second part of this study on the inverse design of metastructure using geometric attention-based deep architecture is a benchmark in the field of inverse design of complex structure whether it is metastructure for the desired mechanical response or any other complex pattern for desired physical properties. The major challenge in the training of an AI-based inverse design model is the availability of a sufficient number of training samples. The higher the complexity of the structure, the higher the number of samples required for the training. Our proposed model uses FEM to learn the dependencies of adjacent connections in the geometrical structure. Therefore, it can easily extract the information embedded in the complex structure and learn the relationship between the input geometry and the output transmissibility. Further, the multiscale geometrical attention mechanism provides the capability to learn the multiple levels of complexity. With the same concept, multiscale feature transformation using the multiscale residual network enables the inverse design model to reconstruct multiple levels of geometry, like a honeycomb and hierarchical honeycomb substructures. Therefore, the proposed framework of the geometrical attention-based feature learning and the inverse design can be useful for any other scientific design problem.
@@
  • 1.
@@
  • 3.
    Creation of novel forward analysis model with a multi-head FEM-inspired spatial attention (FSA) to predict the transmissibility of a metastructure subjected to base excitation. The Multihead FSA efficiently captures the geometrical dependency of vibration mitigation at different scales of structural complexity.
  • 4.
    The newly developed model is highly efficient as it is trained only on 648 training samples. This results in predicting the transmissibility with the mean square error of 0.08 while in similar literature, ML models were generally trained on 5000 to 40000 data sets to achieve a similar level of accuracy.
  • 5.
@@
    Development of a novel MGSA-based inverse design model with 1D spectrum position encoding and a multiscale residual network to produce complex metastructures, given the transmissibility spectrum for a desired frequency bandgap. The proposed inverse design model is found to be much more efficient for the reconstruction of complex metastructure with a similarity score of 0.991.
  • 6.
    Further, we have been able to successfully demonstrate the effectiveness of the proposed inverse design model to create complex multiscale metastructures by providing artificial transmissibility for a desired frequency bandgap. We believe this will facilitate the generation of many customizable complex patterns for mechanical metastructures efficiently.
6 Conclusions
@@
This research study highlights the translational application potential of the vast range of patterns and motifs developed in the field of architecture into metamaterial-based wave attenuation. Thus, the utilitarian values of aesthetic constructs in the field of vibration and sound engineering may open up new systems for wave control, including bandwidth control, sub-wavelength acoustic imaging, and generation of acoustic black holes. Additionally, a systematic investigation process has been created on the basis of root lattice identification, multi-scale lattice formation through a hierarchic system, development of hybrid lattice structure, and the effect of interlacing in the vibration mitigation potential. In general, the simple lattice structures typically act as attenuators, however, this study demonstrates that the interlaced structure offers an enhanced vibration attenuation property depending on the interlace geometry with improved control over the attenuation frequency range. Such effects can be further enhanced in the presence of and through the strategic positioning of the metallic inserts inside the lattice structure, which may result in a significant expansion of the bandwidth. This method has been enhanced with the help of AI-based inverse design models to predict metastructures that exhibit the desired bandgap and the corresponding transmissibility. Although this study concentrates on predominantly honeycomb-shaped interlaced structures, the principle of incorporating mass resonators can be applied to other geometries to manipulate bandgap frequencies. Moreover, the concept may be extended to three-dimensional interlaced structures with diverse lattice and interlace configurations across multiple length scales. Another possibility of future expansion is related to the insertion of smart materials in the lattice structure such that the stiffness and the damping can be controlled over time; thus, transforming it into 4-D metamaterial.
@@ Acknowledgements
Tanuj Gupta: Conceived and expanded the idea jointly, modeled all the structures, and designed and performed numerical simulations. Fabricated the samples by 3D printing and carried out related experiments and material testing, documented the results, reviewed the manuscript, critically examined the results, and checked the data, wrote the manuscript. Arun Kumar Sharma: Concived the idea of AI-based inverse design of metastructures, and implemented a new strategy of feature learning using multiscale GA and the idea of reproducing the geometry with multiple levels of complexity using the multiscale residual network. Code Implementation, training, and validation of the AI models and writing the AI-based inverse design in the manuscript. Critically examined the manuscript. Ankur Dwivedi: Conceived and expanded the idea jointly for the inverse design of metastructures, critically examined the results, checked the data, wrote and reviewed the manuscript. Vivek Gupta: Expanded the idea jointly, reviewed the manuscript, critically examined the results, checked the data, and wrote the manuscript. Subhadeep Sahana: Carried out numerical simulations to generate synthetic data for training the AI model and critically analyzed the paper for logical conclusions. Suryansh Pathak: Conducted the experiments, performed numerical simulations, and checked the data. Ashish Awasthi: Carried out numerical simulations to generate synthetic data for training the AI model. Bishakh Bhattacharya: Conceived the primary idea, reviewed the manuscript, critically examined the results, checked the data, wrote the manuscript, and acquired funding.